What will be the probable value of pressure potential for reverse osmosis if solute potential is -20??

Grade:12th pass

1 Answers

Yuvraj Singh
2 years ago

onΨw=Ψs+ΨρΨw=Ψs+Ψρ
WhereΨw=Ψw=water potential ,Ψs=Ψs=osmotic potential
andΨρ=Ψρ=pressure potential
Ψw=(−20)+(+6)=−14Ψw=(-20)+(+6)=-14"bars"`